# D-LiFT: Improving LLM-based Decompiler Backend via Code Quality-driven Fine-tuning

### v1

@@ 6.2.2 Findings
To better understand D-LiFT’s influence on LLM performance, we conducted an additional analysis by directly comparing the fine-tuned output and the baseline model output. Specifically, we analyze how D-LiFT modifies function performance through two distinct categories: improvements, where fine-tuned models successfully resolve issues present in baseline outputs, and regressions, where previously error-free baseline functions develop new problems following the fine-tuning process. Based on it, we classify these changes into six categories: syntax fixes, semantic fixes, syntax regressions, semantic regressions, readability improvements, and readability regressions. Table IV reports the number of functions falling into each category. These results show that D-LiFT achieves considerable accuracy enhancements for the majority of models, with regression instances remaining at acceptably low levels.
@@
To investigate the root causes of both improvements and regressions, we randomly selected 50 functions from each group (i.e., those that improved and those that regressed) and performed a manual code review. Our findings, summarized in Table V, show that fine-tuning corrects five main categories of errors: (1) Missing instructions, e.g., missing goto labels, variable declarations, and value assignments. (2) Incorrect brackets in conditional expressions, e.g., missing parentheses in if statements or misplacement leading to incorrect pointer dereference. (3) Incorrect variable naming and casting, e.g., using mode_t when only __mode_t is defined. (4) Unwanted instruction insertions, e.g., duplicate goto labels and unwanted function calls. (5) Incorrect literal value. e.g., constant values are inconsistencies with the original values. However, fine-tuning also introduces regressions in three areas: (1) Missing instructions. (2) Incorrect brackets in conditional expressions. (3) Unwanted instruction insertions. We provide more representative examples of the enhancements that D-LiFT enables in Section 6.4.
@@
Summary: When the original decompiled code does not have syntax or semantic error, D-LiFT yields substantial improvement over the baseline, in both the number of functions improved and their average D-Score.
6.2.3 Training Efficiency
Figure 7: This table shows how many functions from the OA dataset scored higher (y-axis) after being processed by the fine-tuned LLMs with different sizes of the training sets (x-axis), compared to the original decompiled code under D-Score.
To determine the optimal training set size, we select our best-performing LLM, i.e., Qwen2.5-Coder, to evaluate LLM performance with training sets of 50, 100, 150, 200, 250, and 300 functions. Figure 7 shows how many functions from the OA dataset scored higher after being processed by the fine-tuned LLM, compared to the original decompiled code under D-Score. As shown, the performance stabilizes once the training set reaches around 200 to 300 functions.
@@ 6.2.3 Training Efficiency
6.3 D-LiFT Performance On OIA
We show the evaluation of D-LiFT on originally inaccurate functions (OIA) in two steps. First, in Section 6.3.1, we use D-Score to compare outputs from the baseline and fine-tuned models. Then, in Section 6.3.2, we analyze the underlying reasons for these observed scores.
6.3.1 Result
Table VI summarizes, for both baseline and fine-tuned models, how many functions achieve a higher D-Score than the original decompiled output and the average score improvement for those functions.
Both the baseline model and the models that are fine-tuned with 300 instances struggle with improving the original inaccurate functions. By taking the best output among all six LLMs and the original decompiled code, only 44 functions showed improvement (average -0.759), while 504 functions remained unimproved (average -2.50).
6.3.2 Findings
To investigate why LLMs perform poorly on the OIA dataset, we randomly select 50 functions, where 25 are successfully improved by any LLMs (including baseline LLMs and fine-tuned LLMs) and 25 are not improved by LLMs, and manually analyze the underlying factors.
Summary: When the original decompiled code contains syntax or semantic errors, neither baseline LLMs nor our fine-tuned models can deliver real improvements, indicating the importance of the decompiler front-end and limitations of current LLMs.
6.4 Case Study
In this section, we present several illustrative cases that demonstrate how D-LiFT effectively improves the quality of decompiled code generated by the LLM without fine-tuning. Specifically, we show how D-LiFT fixes syntax errors introduced by the LLM in Section 6.4.1, fixes semantic errors introduced by the LLM in Section 6.4.2, and improves the readability in Section 6.4.3.
6.4.1 Syntax Error Fixes
@@
Fixing Incorrect Variable Casting and Refactor Loop. Figure 8 illustrates how D-LiFT enhances the output of Qwen2.5-Coder-3B model for the decompiled snippet of the uuid_copy function (util-linux).
@@
if(!tb || !tb->colors_wanted || tb->format != SCOLS_FMT_HUMAN)
@@
if (((*(byte *)(param_1 + 0x140) & 2) == 0) || (*(int *)(param_1 + 0x108)!= 0)) return;
@@
Fixing Missing Parentheses and Consolidate Conditionals. Figure 9 demonstrates another example about how D-LiFT corrects the syntax errors while enhancing readability in the decompiled fputs_color_cell_close (util-linux).
6.4.2 Semantic Error Fixes
@@
Fixing Incorrect Literal Value and Extract Functional Variable. Figure 10 shows how the decompiled code snippet from the fdisk_delete_all_partitions function (util-linux) is improved by correcting the semantic errors while enhancing the readability.
@@
6.4.3 Readability Improvement
Source:
if (!l)
return 0;
for (; *l; l++)
n++;
return n;
Original decompiler:
if (param_1 == (long *)0x0) {
return 0;
if (*param_1 != 0) {
do {
param_1 = param_1 + 1;
iVar1 = iVar1 + 1;
} while (*param_1 != 0);
return iVar1;
return 0;
Baseline model:
if (param_1 == NULL) {
return 0;
if (*param_1!= 0) {
do {
param_1 = param_1 + 1;
iVar1++;
} while (*param_1!= 0);
return iVar1;
return 0;
Fine-tuned model:
if (!param_1) {
return 0;
if (*param_1) {
do {
param_1++;
iVar1++;
} while (*param_1);
return iVar1;
return 0;
Figure 11: This code snippet shows how D-LiFT helps Qwen2.5-Coder-3B enhance its ability to generate more readable code by reducing 29 characters while preserving the accuracy.
Arithmetic Operators & Boolean Expression Simplification. Figure 11 illustrates how D-LiFT enhances readability by simplifying both arithmetic operators and Boolean expressions.
Specifically, the source code, extracted from strv_length function (util-linux), contains 38 non-blank characters. Ghidra’s original decompilation, however, leaves all operators unsimplified, expanding the snippet to 138 non-blank characters. The baseline Qwen2.5-Coder-1.5B model makes marginal improvements, reducing 11 characters in code by simplifying one Boolean expression in lines 6 and 7 of the baseline model output. In contrast, our fine-tuned model systematically rewrites both arithmetic and logical operations, bringing a total of 29 non-blank character reductions while preserving semantics. Consequently, D-LiFT boosts this function’s D-Score from +0.605 to +0.790.
7 Discussion
@@
Applicability. As illustrated in Section 6.3.2, D-LiFT performs poorly when the decompiler front-end fails to produce accurate code. Moreover, because D-LiFT relies on function-level decompiled output, it cannot help if the decompiler is unable to process the binary. Hence, D-LiFT performs poorly in addressing internal decompiler issues, such as bugs that cause decompiler crashes or persistent decompilation challenges, including the identification of function boundaries and indirect calls.
@@
Underlying tools. Since D-LiFT employs D-helix, it inherits the limitations of D-helix, including missing support for floating point instructions, missing support for double pointers (due to the memory model), and missing support for large binaries because of timeouts.
@@
External function calls. For external function calls, D-Score differs from D-helix’s approach of modeling return values as the sum of the least significant bytes of its arguments; Instead, D-Score counts the number of calls. This simplification can be problematic when a function’s return value affects control flow. By default, we assign zero (0) as the return value of every external function call, which may prevent exploration of branches that require nonzero return values. Nevertheless, accurately determining whether an external call yields a meaningful return itself is an open challenge [34]. We leave the more precise modeling of external function calls as future work.
8 Related Work
@@
Regarding LLM-based decompiled-code enhancement, several approaches have been proposed. Besides LLM4Decompiler [55], researchers have also proposed DecGPT [63], which focuses on making decompiled code more recompilable. Nevertheless, this approach does not provide any method to handle the semantic errors, e.g., hallucination errors, introduced by the LLM. To the best of our knowledge, DeGPT [25] is the only existing work that attempts to validate the accuracy of LLM-generated decompiled code. DeGPT introduces MSSC, a static analysis framework that assigns random values to inputs and observes the resulting changes in symbolic values in both the decompiled code and the LLM-generated code, aiming to detect discrepancies and identify inaccuracies. However, this method has notable drawbacks. For instance, MSSC does not recompile the code; it cannot detect syntax errors. Meanwhile, by testing with random inputs, the validation result can be inconsistent across runs.
@@
Regarding the code readability metric, subsequent studies have been proposed based on B&W framework. Specifically, researchers [23, 51, 47, 42, 57] have shifted toward the broader concept of code comprehensibility, which extends readability by also considering elements beyond the code itself, such as associated documentation, during evaluation. However, because these metrics assume the presence of comments and external documentation, features that decompiled code typically lacks, they are not well suited for assessing decompiled output.
9 Conclusion
We design D-LiFT, an automatic decompilation pipeline with an LLM-based back-end that is fine-tuned using reinforcement learning to improve the quality of the decompiled code, adhering to the principle of preserving accuracy while improving readability. We propose D-Score, an integrated scoring mechanism designed specifically for decompilation recovery tasks. We implement D-LiFT based on Ghidra and fine-tune three LLMs, and achieve significant decompiled code improvement for widely used benchmark functions. Compared to the baseline LLMs, on average, our fine-tuned LLMs improve the quality of 55.3% more functions. Moreover, for functions that are accurately decompiled, when choosing the best output among the baseline model, the fine-tuned model (i.e., D-LiFT), and the original decompiled code, we find that on average, 47.3% of the top-scoring functions come from D-LiFT, whereas only 20.9% come from the baseline model.
Ethics Statement

### v2

@@ VI-B2 Findings
To better understand the training part of the D-LiFT’s influence on LLM performance, we conducted an additional analysis by directly comparing the fine-tuned output and the baseline model output. Specifically, we analyze how the training within D-LiFT modifies function performance through two distinct categories: improvements, where fine-tuned models successfully resolve issues present in baseline outputs, and regressions, where previously error-free baseline functions develop new problems following the fine-tuning process. Based on it, we classify these changes into six categories: syntax fixes, semantic fixes, syntax regressions, semantic regressions, readability improvements, and readability regressions. Table˜IV reports the number of functions falling into each category. These results show that LLMs fine-tuned by D-LiFT achieve considerable accuracy enhancements, with regression instances remaining at acceptably low levels.
@@
To investigate the root causes of both improvements and regressions, we randomly selected 50 functions from each group (i.e., those that improved and those that regressed) and performed a manual code review. Our findings, summarized in Table˜V, show that fine-tuning corrects five main categories of errors: (1) Missing instructions, e.g., missing goto labels, variable declarations, and value assignments. (2) Incorrect brackets in conditional expressions, e.g., missing parentheses in if statements or misplacement leading to incorrect pointer dereference. (3) Incorrect variable naming and casting, e.g., using mode_t when only __mode_t is defined. (4) Unwanted instruction insertions, e.g., duplicate goto labels and unwanted function calls. (5) Incorrect literal value. e.g., constant values are inconsistent with the original values. However, fine-tuning also introduces regressions in three areas: (1) Missing instructions. (2) Incorrect brackets in conditional expressions. (3) Unwanted instruction insertions. More representative examples of the enhancements that D-LiFT enables are in Section˜VI-D.
@@
Summary: When the original decompiled code does not have syntax or semantic error, D-LiFT yields substantial improvement, in both the number of functions improved and their average D-Score.
@@
if(!tb $|$ !tb->colors_wanted $|$ tb->format != SCOLS_FMT_HUMAN)
@@
if (((*(byte *)(param_1 + 0x140) & 2) == 0) $|$ (*(int *)(param_1 + 0x108)!= 0)) return;
@@
VI-C D-LiFT Performance on OIA
We show the evaluation of D-LiFT on originally inaccurate functions (OIA) in two steps. In Section˜VI-C1, we compare outputs from the baseline and fine-tuned models. In Section˜VI-C2, we analyze the underlying reasons for these observed scores.
VI-C1 Result
Table˜VI summarizes, for both baseline and fine-tuned models, how many functions achieve a higher D-Score than the original decompiled output and the average score improvement for those functions. As shown, both the baseline model and the fine-tuned models struggle with improving the original inaccurate functions. Moreover, by selecting the best among all six LLMs’ output and the original decompiled code, only 44 functions showed improvement (average -0.759), while 504 functions remained unimproved (average -2.50).
VI-C2 Findings
To investigate why our LLMs perform poorly on the OIA dataset, we randomly select 50 functions, where 25 are successfully improved by any LLMs (including baseline LLMs and fine-tuned LLMs) and 25 are not improved by LLMs, and manually analyze the underlying factors.
Summary: When the original decompiled code contains syntax or semantic errors, neither baseline LLMs nor our fine-tuned LLMs can deliver real improvements, indicating the importance of the decompiler front-end and the limitation of current LLMs in fixing those errors introduced by the decompiler.
VI-D Case Study
In this section, we present several illustrative cases that demonstrate how D-LiFT effectively improves the readability of decompiled code while preserving the accuracy. Specifically, we show how D-LiFT fixes syntax errors introduced by the LLM in Section˜VI-D1, fixes semantic errors introduced by the LLM in Section˜VI-D2
VI-D1 Syntax Error Fixes
Fixing Incorrect Variable Casting and Refactor Loop. Figure˜8 illustrates how D-LiFT enhances the output of Qwen2.5-Coder-3B model for the decompiled snippet of the uuid_copy function (util-linux).
Fixing Missing Parentheses and Consolidating Conditionals. Figure˜9 demonstrates another example about how D-LiFT corrects the syntax errors while enhancing readability in the decompiled fputs_color_cell_close (util-linux).
VI-D2 Semantic Error Fixes
Fixing Incorrect Literal Value and Extracting Functional Variable. Figure˜10 shows how the decompiled code snippet from the fdisk_delete_all_partitions (util-linux) is improved by fixing the semantic errors while enhancing the readability.
@@ VI-D2 Semantic Error Fixes
VII Discussion
@@
Applicability. As illustrated in Section˜VI-C2, D-LiFT performs poorly when the decompiler front-end fails to produce accurate code. Moreover, because D-LiFT relies on function-level decompiled output, it cannot help if the decompiler is unable to process the binary. Hence, D-LiFT performs poorly in addressing internal decompiler issues, such as bugs that cause decompiler crashes or persistent decompilation challenges, such as function boundaries identification, type recovery, or indirect calls. We point out that decompilation is only one of multiple aspects of binary reverse engineering, focusing on presenting users with more readable source code of the subject binary. The other aspects of reverse engineering are better supported by other specialized reverse engineering tools, such as those for type recovery [35, 12], control flow graph reconstruction [7], and function boundary recovery [65], which complement decompilation.
@@
Underlying tools. Since D-LiFT employs D-helix, it inherits the limitations of D-helix, including missing support for floating-point instructions, double pointers (due to the memory model), and large binaries because of timeouts.
@@
External function calls. For external function calls, D-Score differs from D-helix’s approach of modeling return values as the sum of the least significant bytes of its arguments; Instead, D-Score counts the number of calls. This simplification can be problematic when a function’s return value affects control flow. By default, we assign a constant value (e.g., 0) as the return value of every external function call, which may prevent exploration of branches. Nevertheless, accurately determining whether an external call yields a meaningful return itself is an open challenge [34]. We leave the more precise modeling of external function calls as future work.
VIII Related Work
@@
Regarding LLM-based decompiled-code enhancement, several approaches have been proposed. Besides LLM4Decompiler [56], researchers have also proposed DecGPT [64], focusing on making decompiled code more recompilable. Nevertheless, this approach does not handle the semantic errors, e.g., hallucination errors, introduced by the LLM. To the best of our knowledge, DeGPT [25] is the only existing work that attempts to validate the accuracy of LLM-generated decompiled code. DeGPT introduces MSSC, a static analysis framework that assigns random values to inputs and observes the resulting changes in symbolic values in both the decompiled code and the LLM-generated code, aiming to detect discrepancies and identify inaccuracies. However, this method has notable drawbacks. For instance, MSSC does not recompile the code; it cannot detect syntax errors. Meanwhile, by testing with random inputs, the validation result can be inconsistent across runs.
@@
Regarding the code readability metric, subsequent studies have been proposed based on the B&W framework. Specifically, researchers [23, 52, 48, 42, 58] have shifted toward the broader concept of code comprehensibility, which extends readability by also considering elements beyond the code itself, such as associated documentation, during evaluation. However, because these metrics assume the presence of comments and external documentation, features that decompiled code typically lacks, they are not well suited for assessing decompiled output.
IX Conclusion
We design D-LiFT, an enhanced decompiler-LLM pipeline with a fine-tuned LLM using code quality-aware RL to improve the quality of the decompiled code, adhering to the principle of preserving accuracy while improving readability. We propose D-Score, an integrated scoring mechanism designed specifically for decompilation recovery tasks. We implement D-LiFT based on Ghidra and fine-tune three LLMs, and achieve significant decompiled code improvement for widely used benchmark functions. Compared to the baseline LLMs, on average, our fine-tuned LLMs improve the quality of 55.3% more functions. Overall, D-LiFT generates 68.2% better quality functions than the native decompiler, where 47.3% of functions come from the D-Score-driven fine-tuned model and only 20.9% from the baseline model.
Ethics Considerations